OpenStudy (anonymous):

Mr Simpson builds a wire fence around his field. The field is 350 metres wide and 450 metres long. He has a wooden gates on two of the sides, each of which is 4 metres wide. What is the total length of wire fencing?

OpenStudy (anonymous):

2(350+450)-2*4=800-16=784

OpenStudy (anonymous):

=2*(800)-16 =1600-16 =1584

OpenStudy (anonymous):

The perimeter without gates is 2*(350+450) = 1600 There are 2 gates on each of the sides gives 4 gates in total 1600- 4*4 = 1584

OpenStudy (anonymous):

Although this isn't clear as it could be 2 gates or 4 gates gates

OpenStudy (anonymous):

ok thankyou